A bag of peanuts is worth 30 cents less than a bag of cashews. Equal amounts of peanuts and cashews are mixed to make 40 bags of a mixture that sells for $1.05 per bag. How much would a bag of cashews be worth?
Answer by ptaylor(2198)


We'll deal in pennies, OK?
Let x=value of a bag of Cashews
Then x-30=value of a bag of peanuts
We know that the value of the cashews(20*x) and peanuts(20(x-30)) before they are mixed together has to equal the value of the final mixture(40*105), so our equation to solve is:

20x+20(x-30)=40*105 get rid of parens
20x+20x-600=4200 add 600 to each side
20x+20x-600+600=4200+600 collect like terms
40x=4800 divide each side by 40
x=120 or $1.20 value of a bag of cashews
CK
20(120)+20(90)=40*105
2400+1800=4200
4200=4200
Hope this helps---ptaylor